Suppose you do 5.7 J of work on a balloon by compressing it. In the process, the balloon loses 3.3 J of heat. What is the change in internal energy of the balloon?
1
Expert's answer
2020-04-22T10:54:49-0400

∆U = Q + W

Q = - (heat lost) = - 3.3 J

W = + (work done on system) = + 5.7 J

∆U = - 3.3 + 5.7 = + 2.4 J